Saved Lost
First
Class
Men 57 116
Women 140 4
Children 6 1
Second Class Men 14 135
Women 80 24
Children 24 0
Third
Class
Men 72 385
Women 74 88
Children 32 57
Crew & Employees Women 20 0
Men 192 685

Given a survivor of the Titanic is a man. what is the probability, rounded to the nearest whole percent, the survivor was a second class passenger (based on the data above)
Answer #1

Total Survivor Man = 57 + 14 + 72 + 192 = 335

P(Survivor Second Class/Survivor Man) = P(Survivor Second class AND Survivor Man)/P(Survivor Man)

                                                       = 14/335 =0.0418

So,

Answer is:

0.0418